If X-3Y=8 and Y=2x+1; X=?
Answer by RAY100(1637)

(1) x-3y=8
(2) y=2x+1
.
subst (2) into (1)
.
x-3(2x+1) =8
x-6x -3 =8
-5x=11
x=-11/5
.
y= 2(-11/5) +1
y=-22/5 +5/5
y=-17/5
.
check,,,(1) (-11/5) -3(-17/5) =8
-11/5 + 51/5 = 40/5 =8,,,,ok
.
(2) -17/5 =2(-11/5) +1
-17/5 = -22/5 +5/5 = -17/5,,,,ok